could you please help me with factor 5xsquare+17x+6
Found 2 solutions by chelcilovesmj1, Alan3354:
Answer by chelcilovesmj1(3)


5x^2 + 17x + 6
6 * 5 = 30
30 / 2 = 15
15 + 2 = 17
now we can break 17 down into 15 and 2
5x^2 + 15x + 2x + 6
factor out a 5x and 2x we get
5x(x+3) + 2(x+3)
so we have (5x+2) and (x+3)
Final answer (5x+2)(x+3)
hope this helps :)


Answer by Alan3354(69443)

could you please help me with factor 5xsquare+17x+6
-------------
It will be (x + a)*(5x + b)
a & b are positive, and their product is 6.
Only 3 possibilities for a & b.